# A Greedy Quantum Route-Generation Algorithm

### v1

@@ IV-A Greedy Quantum Route-Finding Algorithm
7:         Compute $|<Z_{i}>|$ over all $M$ samples
@@
9:     $\bar{X}^{l},\leftarrow$ Select $\theta$ of $X^{l}$ with highest $|<Z_{i}>|$
@@
Step 8 computes a statistic used in step 9 to determine which variables will participate in the DAG for finding paths. In our algorithm, we used one-body expectation values $|<Z_{i}>|$, although two-body expectation values may be used [12]. Parameter $\theta\in(0,1)$ is given by the user to determine the variables in $X^{l}$ to select at each iteration. $\theta$ may be given as a portion of the variables to select, or as a threshold for expectation values. In our running example and our analysis, we specify $\theta$ as a portion, although the results still hold for $\theta$ as a threshold. For step 10, any sub-routine may be used that guarantees that no 2 paths returned visit same customer. This ensures that the paths are edge-and-vertex-disjoint. Two examples are: returning the longest path, or returning multiple paths. Multiple paths may be returned by an iterative procedure: 1. Select the longest path in the DAG. 2. Remove all nodes (and their adjacent edges) in the DAG with the same customers $i_{k}$ (excluding depots $0,N$) along the currently-selected longest path, 3. Remove the nodes and arcs along the currently-selected longest path from the DAG. Repeat until the DAG is empty. In the simulation results, we returned multiple paths using this procedure. Heuristics such as returning a single longest path are more conservative, and may lead to longer computational time since fewer variables would be fixed at each iteration.
IV-B DAG Representation
We now describe the construction of the DAG. Any set of active variables $X^{l}$ may be represented as follows: A single node is created for each unique tuple $(i,s)$, where $i\in\{0,W,N\}$ and $s\in T$. A directed arc is present from $(i,s)$ to $(j,t)$ for each variable $x_{i,s,j,t}\in X^{l}$. By constraints (3g) and (3h), after the pre-processing described in section III-A variables only remain active if $s<t$. There cannot exist a cycle, otherwise it would imply that there exists a variable $x_{i,s,j,t}$ such that $s\geq t$. DAG($X^{0}$) for our running example is shown in Fig. 1.
IV-C Pruning
The subroutine Prune($S,X,Q$) is shown in Algorithm 2. Notice that the constraints (3b) and (3c) are separable by customer $j$, and tuples $(i,s)$ respectively. So we may prune variables and remove constraints independently for each tuple $(i,s)$, along a path in steps 6, 8, and 11 in Algorithm 2. If a path is found that shares an endpoint (either start or end) of a path in $S$, the paths are concatenated by Concat and the appropriate constraints are removed, as shown in Algorithm 3.
@@ IV-C Pruning
1:$S$, set of paths found previously, $Q$, new set of paths found.$X$, current set of active variables.
@@
6:         $X^{\prime}\leftarrow$ Apply pruning rules in section IV-C1.
@@
8:              QUBO $\leftarrow$ remove constraint (3c,$i_{k},s_{k}$)
@@
11:              QUBO $\leftarrow$ remove constraint (3b,$i_{k}$)
@@
Figure 1: Initial set of variables $X^{0}$ represented as a DAG (denoted DAG($X^{0}$)), continued from example in Tab. I, labeled with possible one-body expectation values from step 8 in Algorithm 1. Variables selected in step 9 in Algorithm 1, using $\theta=0.5$ are marked with solid arrows. Paths found in step 10 in Algorithm 1 are highlighted in blue.
Figure 2: Pruning example for $\theta=0.9$. Arcs in paths found in step 10 of Algorithm 1 are highlighted blue, and variables along the path are set to $1$ and pruned. Arcs highlighted red are set to 00 and pruned. Paths found are $P_{1}=\{(1,1),(N,2)\}$, $P_{2}=\{(0,0),(2,3),(N,3)\}$.
TABLE III: Pruning Rules Applied to Fig. 2
@@
Figure 3: Subsequent iteration for pruning example with $\theta=0.9$. (Expectation values omitted to avoid clutter). After pruning in previous iteration, only variable $x_{0,0,1,1}$ remains active. Path $P_{3}=\{(0,0),(1,1)\}$ is found in step 10 of Algorithm 1, and joined with path $P_{1}$ in step 9 of Algorithm 3.
IV-C1 Pruning Rules
Throughout this section, the phrase ”prune” means the variable is set to the value specified and removed from the set of active variables $X^{l}$. For customer $i_{k}$ of tuple $(i_{k},s_{k})$, variables involving customer $i_{k}$ may be classified as:
  1. 1.
    $i=i_{k}$: ”outgoing” variables.
  2. 2.
    $j=i_{k}$: ”incoming” variables.
Index $k\in 0,...|P|$ along a path $P$ may be classified
  1. 1.
    $k\notin\{0,|P|\}$: ”interior” indices.
  2. 2.
    $k\in\{0,|P|\}$: ”exterior” indices.
@@ IV-C1 Pruning Rules
    ”outgoing” variables are set to 00 and pruned if either:
@@
    ”incoming” variables are set to 00 and pruned if either:
@@
      $i_{k}=0$: Only prune and set $=1$ if $x_{i_{k},s_{k},i_{k+1},s_{k+1}}$ is Along path. (Variables incoming to depot were pruned by (3d)).
@@
      $i_{k}\neq 0$: Prune using rules ”interior” 1 and 2a.
@@
      $i_{k}=N$: Prune using ”interior-incoming” 2a. (We do not prune according to rules 2b because incoming arcs to final depot $N$ at different times is allowed, as shown in Fig. 1).
@@ V Analysis
    $i^{P}_{k}$ was in the interior of either path $R$ or $R^{\prime}$. So it was either pruned to $1$ by the ”Along path” rules, or pruned to 00 by the ”interior” rules in section IV-C.
@@
    $i^{P}_{k}$ was in the exterior of of either path $R$ or $R^{\prime}$. The incoming variables to $i^{P}_{k}$ were pruned in iteration $l^{R}$ by ”exterior” rule 2b and the outgoing variables were pruned in iteration $l^{R^{\prime}}$ by ”exterior” rule 1b.
@@
For every customer $i^{\prime}$, $\exists$ iteration $l^{\prime}$ in which the paths $Q^{l^{\prime}}$ returned in step 10 of Algorithm 1 will contain a path with a tuple $(i^{\prime},s)$ for some $s$.
@@
    Annealer returns 0 for all variables, for all samples. Then no variables are pruned, and the while loop resumes.
@@
      $\bar{X}^{l}$ does not contain a variable $x_{i,s,j,t}$ with either $i^{\prime}=i$ or $i^{\prime}=j$. Variables containing customers along paths returned by step 10 will be pruned according to the pruning rules in section IV-C1, and the while loop resumes its next iteration with $X^{l+1}\subset X^{l}$.
@@
In case 2(b)ii the variables containing customers along paths returned by step 10 will be pruned according to the pruning rules in section IV-C1, and the while loop resumes its next iteration with $X^{l+1}\subset X^{l}$. If in case 2(b)i. QED. If the annealing time $A$ is sufficiently large, [17] eventually we will exit case 1. If we continue to be in cases 2a or 2(b)ii above, then paths returned in step 10 in Algorithm 1 do not contain tuple $(i^{\prime},s)$, and in the next iteration $X^{l+1}\subset X^{l}$. By Lemma 1, all variables in the interior of previously-found paths $S$ will be pruned. At some iteration $l^{\prime}$ the only variables not pruned are those containing tuple $(i^{\prime},s)$, in which case they must appear in the paths returned in step 10. ∎
@@
If $k=0$ then all outgoing variables and incoming variables to $i^{R}_{0}$ with $t\neq s^{R}_{0}$ have been pruned in ”exterior” rule 1b. So, the only active variables involving $i^{R}_{0}$ have $(j,t)=(i^{R}_{0},s^{R}_{0})$. By Lemma 2, in iteration $l^{\prime}$ path $R^{\prime}\in Q^{l^{\prime}}$ will contain $(i^{R}_{0},s^{R}_{0})$. This can only happen if $i^{R^{\prime}}_{|R^{\prime}|}=i^{R}_{0}$. In step 9 in Algorithm 3 these paths are unioned, and the resulting path $P=R^{{}^{\prime}\,{}^{\frown}}R$ will have $i^{R}_{0}$ in its interior.
The proof for $k=|R|$ is the similar, except $i^{R^{\prime}}_{0}=i^{R}_{|R|}$, and the resulting path is $P=R^{\frown}R^{\prime}$. ∎
@@
Algorithm 1 converges to a feasible solution to Problem (3).
@@
Observe that all variables contain a customer by pre-processing constraint (3l). By Lemma 3 at some iteration, any customer $i$ will be in the interior of some path $P\in S$. By the pruning rules, when a customer is in the interior of a path $P\in S$ we have set exactly two variables involving customer $i$ equal to $1$: one incoming, and one outgoing. All other variables involving customer $i$ have been pruned and set to 00. Thus constraint (3b) is satisfied. By the same argument, if $(i,s)$ is along a path in $S$ returned by Algorithm 1 then exactly one variable on each side of constraint (3c,$i,s$) will be $1$ all others have been pruned by rules ”interior” 1 and 2. If $(i,s)$ is not along a path, then constraints (3c, $i,s$) will become $0=0$ and will be satisfied again by rules ”interior” 1 and 2. Since constraints (3b) and (3c) are satisfied the paths $S$ returned by Algorithm 1 are a feasible solution to Problem (3). ∎
VI Computational Results
VI-A Solvers
@@
We benchmark with both classical and quantum annealing approaches that attempt to find a feasible solution to the FSVRPTW. In the absence of obvious heuristics for restoring feasibility, the standard approach is to run quantum annealer to obtain a distribution of samples, and select the feasible sample with the lowest energy. We give a brief description of each method used:Simulated annealing (SAS) is a classical metaheuristic algorithm that mimicks the physical process of annealing. Its algorithmic parameters include an annealing schedule. We use the SimulatedAnnealingSampler (SAS) provided by D-Wave, and we use the default settings of a geometric schedule [18]. LeapHybridCQMSampler (CQM) is a hybrid classical/quantum annealer that is specifically designed to solve problems with constraints. Its algorithmic parameters include num_reads, which is the number of samples returned by the annealer. The more samples, the greater chance of finding a solution with low energy. We used D-Wave’s CQM version 1p solver, and set num_reads=1000. In our greedy algorithm we used CQM for solving the QUBO($X$) in step 1, with the default parameters, and the default penalty values. To evaluate the performance of our proposed algorithm for practical-sized problems we used Solomon’s benchmark instances [19], briefly described as graphs with nodes that must be serviced with a time window, placed according to a random distribution. We used instance types $R101$/$R201$ which indicates narrow/wide time windows resp. We modify an instance to suit our problem as follows: for a given number of customers, randomly sample $N$ customers. For each value of $N$, we construct 10 random examples independently, and aggregate their statistics.
@@
From Fig. 4(a), and 4(b) we can see that our greedy approach enjoys a small relative optimality gap as compared to SAS and CQM, in particular for large problem sizes $N>15$. This difference is even more pronounced in the R201 examples. For $N=25$, SAS returned feasible solutions for only 2 of the 10 random examples, and for $N=50$ it returned no feasible solutions. Statistics are aggregated only over feasible solutions. Fig. 5(a), and 5(b) show the relative time difference between the greedy algorithm, and CQM/SAS. Positive values indicate greedy algorithm is faster. For large problem instances greedy is always faster than SAS, although CQM is faster by a small margin for $N>15$.
@@
Figure 4: Average relative optimality gap for SAS, CQM, and the proposed greedy algorithm, averaged over 10 examples with standard deviation. Relative optimal value is computed as $\frac{C_{A}-C_{opt}}{C_{opt}}$ where $C_{A}$ is the objective value found by the annealing method, and $C_{opt}$ is the optimal value found by Gurobi version 10.0.3. Statistics are computed only over feasible solutions. Data not shown for SAS, $N=50$, because no feasible solutions were found.
@@
Figure 5: Average relative time difference of greedy as compared to SAS and CQM averaged over 10 examples. Relative time difference is computed as $\frac{t_{A}-t_{greedy}}{t_{A}}$ where $t_{A}$ is the time taken by the annealing method, and $t_{greedy}$ is the time taken by our proposed greedy algorithm. Statistics are computed only over feasible solutions. Data not shown for SAS, $N=50$, because no feasible solutions were found.
TABLE IV: QPU Results
$N$ DWaveSampler ($A=50\mu s$, $M=2000$) % Valid DWaveSampler ($A=50\mu s$, $M=2000$) with Pause % Valid $t_{\text{greedy}}$ $\alpha$ Avg. # logical Avg. # physical
5 20 10 0.08 10 44 415
6 0 0 0.074 17 84 601
7 0 0 0.08 33 109 890
8 0 0 0.09 0 121 955
9 0 0 0.09 0 127 1021
VII QPU Results
To evaluate its robustness to noise, our algorithm was also tested on DWave Advantage 4.1. For a fair benchmark, we also tried tuning the parameters of the DWaveSampler. An annealing pause was used by setting it to: [0.0, 0.0], [10.0, 0.4], [50.0, 0.4], [120, 1.0]. Annealing pauses are showng to perform well for scheduling problems [20]. Finally, to increase the chance of the DWaveSampler finding a feasible solution, both the annealing time, and the number of samples were adjusted to $50\mu s$ and $M=2000$, respectively. (This is the largest these values can be increased before the DWave’s QPU access time is exceeded [21]). In our greedy algorithm, we left the annealing parameters in DWaveSampler to their default values.
Our experiments revealed that even after enhancing the DWaveSampler with these features, it fails to find any feasible solutions in any of the 10 examples for problem sizes greater than N=5 customers. The greedy algorithm, when using the default DWaveSampler parameters, still maintains an average relative optimality gap below 33%. The total time of Algorithm 1, and the average relative optimality gap are labeled $t_{\text{greedy}}$, and $\alpha$, respectively, Table IV.
To show the scaling of this formulation on the binary variables, the average number of logical qubits and physical qubits are also reported. When the number of customers exceeds 7, the rate at which the number of logical qubits increases more slowly. This can be explained by the fact that only one discretization value within each time window is needed. If there is already one time discretization within a time window, then we avoid creating a new one. As a timetable includes more customers, more of the time windows overlap.
Conclusions and Future Work

### v2

@@ 4.1 Greedy Quantum Route-Generation Algorithm
7:         Compute $\frac{\langle Z_{i}\rangle+1}{2}$ over all $M$ samples
@@
9:     $\bar{X}^{l},\leftarrow$ Select $\theta$ fraction of $X^{l}$ with highest $\frac{\langle Z_{i}\rangle+1}{2}$
@@
Step 8 computes a statistic used in step 9 to determine which variables will participate in the DAG for finding paths. In our algorithm, we used one-body expectation values, although two-body expectation values may be used [5]. Parameter $\theta\in(0,1)$ is given by the user to determine the variables in $X^{l}$ to select at each iteration. $\theta$ may be provided as a portion of the variables to choose or as a threshold for expectation values. In our running example and our analysis, we specify $\theta$ as a portion. For step 10, any sub-routine may be used that guarantees that no two paths returned visit the same customer. Two examples are: returning the single longest path or returning multiple paths. An iterative procedure may return multiple paths: 1. Select the longest path in the DAG. 2. Remove all nodes (and their adjacent edges) in the DAG involving any of the customers along the currently selected longest path (excluding depots $0,N$), 3. Remove the nodes and arcs along the currently selected longest path from the DAG. 4. Repeat until the DAG is empty. In the simulation results, we returned multiple paths using this procedure. Heuristics such as returning a single longest path are more conservative and allow more flexibility for the algorithm to adapt in consecutive iterations but may lead to longer computation time.
@@
1:$S$, set of paths found previously, $Q$, new set of paths found, $X$, current set of active variables.
@@
6:         $X^{\prime}\leftarrow$ Apply pruning rules in §4.3.1.
@@
8:              QUBO $\leftarrow$ remove constraint (4c,$i_{k},s_{k}$)
@@
11:              QUBO $\leftarrow$ remove constraint (4b,$i_{k}$)
@@
4.2 DAG Representation
Any set of active variables $X^{l}$ may be represented as follows: A single node is created for each unique tuple $(i,s)$, where $i\in\{0,W,N\}$ and $s\in T$. A directed arc is present from $(i,s)$ to $(j,t)$ for each variable $x_{i,s,j,t}\in X^{l}$. By constraints (4g) and (4h), after the pre-processing described in §3.1 variables only remain active if $s<t$. There cannot exist a cycle; otherwise, it would imply that there exists a variable $x_{i,s,j,t}$ such that $s\geq t$. DAG($X^{0}$) for our running example is shown in Fig. 1.
4.3 Pruning
The subroutine Prune($S,X,Q$) is shown in Alg. 2. Notice that the constraints (4b) and (4c) are separable by customer $j$, and tuples $(i,s)$ respectively. So we may prune variables and remove constraints independently for each tuple $(i,s)$, along a path in steps 6, 8, and 11 in Alg. 2. If a path is found that shares an endpoint (either start or end) of a path in $S$, the paths are concatenated by Concat and the appropriate constraints are removed, as shown in Alg. 3.
4.3.1 Pruning Rules
Throughout this section, the phrase “prune” means the variable is set to the value specified and removed from the set of active variables $X^{l}$. For customer $i_{k}$ of tuple $(i_{k},s_{k})$, variables involving customer $i_{k}$ may be classified as:
  1. 1.
    $i=i_{k}$: "outgoing" variables.
  2. 2.
    $j=i_{k}$: "incoming" variables.
Index $k\in 0,...|P|$ along a path $P$ may be classified
  1. 1.
    $k\notin\{0,|P|\}$: “interior” indices.
  2. 2.
    $k\in\{0,|P|\}$: “exterior” indices.
@@ 4.3.1 Pruning Rules
(a) Arcs in paths found in step 10 of Algorithm 1 are highlighted blue, and variables along the path are set to $1$ and pruned. Arcs highlighted in red are set to 00 and pruned. Paths found are $P_{1}=\{(1,1),(N,2)\}$, $P_{2}=\{(0,0),(2,3),(N,3)\}$.
(b) After pruning in the previous iteration, only variable $x_{0,0,1,1}$ remains active. Path $P_{3}=\{(0,0),(1,1)\}$ is found in step 10 of Alg. 1, and joined with path $P_{1}$ in step 9 of Alg. 3.
Figure 2: Pruning example for $\theta$=0.9, and subsequent iteration of Alg. 1
Table 2: Pruning Rules Applied to Fig. 2(a)
@@
    “outgoing” variables are set to 00 and pruned if either:
@@
    “incoming” variables are set to 00 and pruned if either:
@@
      $i_{k}=0$: Only prune and set $=1$ if $x_{i_{k},s_{k},i_{k+1},s_{k+1}}$ is Along path. (Variables incoming to the depot were pruned by (4d)).
@@
      $i_{k}\neq 0$: Prune using rules “interior” 1 and 2a.
@@
      $i_{k}=N$: Prune using “interior-incoming” 2a. (We do not prune according to rules 2b because incoming arcs to final depot $N$ at different times are allowed, as shown in Fig. 1).
@@ 5 Analysis
    $i^{P}_{k}$ was in the interior of either path $R$ or $R^{\prime}$. So it was either pruned to $1$ by the “Along path” rules or pruned to 00 by the “interior” rules in §4.3.
@@
    $i^{P}_{k}$ was in the exterior of of either path $R$ or $R^{\prime}$. The incoming variables to $i^{P}_{k}$ were pruned in iteration $l^{R}$ by “exterior” rule 2b and the outgoing variables were pruned in iteration $l^{R^{\prime}}$ by “exterior” rule 1b.
@@
For every customer $i^{\prime}$, $\exists$ iteration $l^{\prime}$ in which the paths $Q^{l^{\prime}}$ returned in step 10 of Alg. 1 will contain a path with a tuple $(i^{\prime},s)$ for some $s$.
@@
    Annealer returns 0 for all variables, for all samples. Then, no variables are pruned, and the while loop resumes.
@@
      $\bar{X}^{l}$ does not contain a variable $x_{i,s,j,t}$ with either $i^{\prime}=i$ or $i^{\prime}=j$. Variables containing customers along paths returned by step 10 will be pruned according to the pruning rules in §4.3.1, and the while loop resumes its next iteration with $X^{l+1}\subset X^{l}$.
@@
In case 2(b)ii, the variables containing customers along paths returned by step 10 will be pruned according to the pruning rules in §4.3.1, and the while loop resumes its next iteration with $X^{l+1}\subset X^{l}$. If in case 2(b)i. QED. If the annealing time $A$ is sufficiently large, [2] eventually, we will exit case 1. If we continue to be in cases 2a or 2(b)ii above, then paths returned in step 10 in Alg. 1 do not contain tuple $(i^{\prime},s)$, and in the next iteration $X^{l+1}\subset X^{l}$. By Lemma 1, all variables in the interior of previously found paths $S$ will be pruned. At some iteration $l^{\prime}$, the only variables not pruned are those containing tuple $(i^{\prime},s)$, in which case they must appear in the paths returned in step 10. ∎
@@
If $k=0$, then all outgoing variables and incoming variables to $i^{R}_{0}$ with $t\neq s^{R}_{0}$ have been pruned in “exterior” rule 1b. So, the only active variables involving $i^{R}_{0}$ have $(j,t)=(i^{R}_{0},s^{R}_{0})$. By Lemma 2, in iteration $l^{\prime}$ path $R^{\prime}\in Q^{l^{\prime}}$ will contain $(i^{R}_{0},s^{R}_{0})$. This can only happen if $i^{R^{\prime}}_{|R^{\prime}|}=i^{R}_{0}$. In step 9 in Alg. 3, these paths are unioned, and the resulting path $P=R^{{}^{\prime}\,{}^{\frown}}R$ will have $i^{R}_{0}$ in its interior. The proof for $k=|R|$ is the similar, except $i^{R^{\prime}}_{0}=i^{R}_{|R|}$, and the resulting path is $P=R^{\frown}R^{\prime}$. ∎
@@
Alg. 1 converges to a feasible solution to Problem (4).
@@
Observe that all variables contain a customer by pre-processing constraint (4l). By Lemma 3 at some iteration, any customer $i$ will be in the interior of some path $P\in S$. By the pruning rules, when a customer is in the interior of a path $P\in S$, we have set exactly two variables involving customer $i$ equal to $1$: one incoming and one outgoing. All other variables involving customer $i$ have been pruned and set to 00. Thus, constraint (4b) is satisfied. By the same argument, if $(i,s)$ is along a path in $S$ returned by Alg. 1, then exactly one variable on each side of constraint (4c,$i,s$) will be $1$ all others have been pruned by rules “interior” 1 and 2. If $(i,s)$ is not along a path, then constraints (4c, $i,s$) will become $0=0$ and will be satisfied again by rules “interior” 1 and 2. Since constraints (4b) and (4c) are satisfied, the paths $S$ returned by Alg. 1 are a feasible solution to Problem (4). ∎
Lemma 5.
If in Alg. 1 step 5 is solved to optimality, and multiple longest paths are returned using the procedure described in section 4.1 then Alg. 1 converges to the optimal solution to Problem (4) in a single iteration.
Proof.
We give the proof for a sub-routine returning multiple longest paths as described in section 4.1. Observe that in the first iteration, step 5 solves Problem (4) with all variables active, so it is equivalent to solving Problem (4). If step 5 is solved to optimality, then it returns a single sample, so the one-body expectation values are all $1$. If $\theta$ is taken to be a threshold strictly between 00 and $1$, then $\bar{X}^{l}$ will only contain variables assigned a value of $1$ by the exact solver. Notice that by returning multiple longest paths as described in §4.1, the collection of paths in $Q^{l}$ will contain all customers. Observe that step 6 will prune the variables in DAG($\bar{X}^{l}$) to a value of $1$, and any variables not in DAG($\bar{X}^{l}$) to a value of 00, which is the same as the solution returned by the exact solver. Since $Q^{l}$ contains paths with all customers, all variables will be pruned, so Alg. 1 will terminate. From Lemma 4, the result will be feasible. Since an exact solver was used, it is also optimal. ∎
Table 3: Notation and Solver descriptions
Notation Description
greedy+QPU Alg. 1 with Step 5 solved using Advantage2
greedy+CQM Alg. 1 with Step 5 solved using LeapHybridCQM
greedy+SAS Alg. 1 with Step 5 solved using SimulatedAnnealingSampler
greedy+Exact Alg. 1 with Step 5 solved using Gurobi
QPU Problem (4) solved using Advantage2
CQM Problem (4) solved using LeapHybridCQM
SAS Problem (4) solved using SimulatedAnnealingSampler
6 Computational Results
@@
We benchmark with both classical and quantum annealing approaches for solving Problem 4. In Algorithm 1, an annealer or a MIP solver may be used in step 5. The solvers described in section 2 were used to either solve Problem 4 or were used to solve step 5 in Algorithm 1. The notation for each solution method is in Tab. 3. For the solution methods QPU, CQM, and SAS, a feasible solution was obtained by sorting samples by energy and selecting the first feasible solution. To evaluate the performance of our proposed algorithm for practical-sized problems, Solomon’s benchmark instances were used [26], briefly described as graphs with nodes that must be serviced within a time window, placed according to a distribution. Instance types $R101$/$R201$ indicate narrow/wide time windows resp. We modify an instance to suit our problem as follows: Randomly sample $N$ customers. For each value of $N$, construct ten random examples independently and aggregate their statistics.
6.1 Results - greedy+CQM, SAS, and CQM
@@
From Fig. 3(a), and 3(b), we can see that the greedy+CQM approach enjoys a small relative optimality gap as compared to SAS and CQM, in particular for large problem sizes $N$>15. This difference is even more pronounced in the $R$201 examples. For $N$=25, SAS returned feasible solutions for only 2 of the ten random examples, and $N$=50, it returned no feasible solutions. Statistics are aggregated only over feasible solutions. Fig. 4(a), and 4(b) show the relative time difference between the greedy algorithm and CQM/SAS. Positive values indicate that the greedy algorithm is faster. For large problem instances, greedy is always faster than SAS. And CQM is sometimes faster by a small margin for $N$>10.
Table 4: greedy+QPU Results: $t$ is absolute time (sec), $\alpha$ is rel. optimality gap, $X$ is the number of logical/physical qubits.
$N$ $t_{R101}$ $\alpha_{R101}$ $X_{log,R101}$ $X_{phys,R101}$ $t_{R201}$ $\alpha_{R201}$ $X_{log,R201}$ $X_{phys,R201}$
5 0.08 0 44 415 0.071 0 24 356
6 0.074 0 84 601 0.075 0 78 585
7 0.08 10 109 890 0.08 10 101 792
8 0.09 17 121 955 0.11 13 109 834
9 0.09 33 127 1021 0.09 27 128 998
@@
Figure 3: Average relative optimality gap for SAS, and CQM and greedy+CQM, averaged over 10 examples with standard deviation. The relative optimal value is computed as $\frac{C_{A}-C_{opt}}{C_{opt}}$ where $C_{A}$ is the objective value found by the annealing method, and $C_{opt}$ is the optimal value found by Gurobi. Statistics are computed only over feasible solutions. Data not shown for SAS, $N$=50 - no feasible solutions found.
@@
Figure 4: Relative time difference between greedy+CQM and SAS /CQM averaged over ten examples. Rel. time difference is computed as $\frac{t_{A}-t_{greedy}}{t_{A}}$ where $t_{A}$ is the time of SAS/CQM, and $t_{greedy}$ is time taken of greedy+CQM. Statistics are computed only over feasible solutions. Data not shown for SAS, $N$=50, because no feasible solutions found.
Table 5: Results for greedy+QPU for R101 and R201; $\alpha$ is % rel. optimality gap, $v$ is % valid solutions, $t$ is embedding time
$N$ $\alpha_{R101}$ $v_{greedy,R101}$ $t_{emb,R101}$ $\alpha_{R201}$ $v_{greedy,R201}$ $t_{emb,R201}$
5 0.0 $\pm$ 0.0 100 0.09 $\pm$ 0.01 0.0 $\pm$ 0.0 100 2.31 $\pm$ 1.40
6 10.0 $\pm$ 2.1 100 2.44 $\pm$ 1.5 8.0 $\pm$ 4.1 100 0.16 $\pm$ 0.04
7 2.5 $\pm$ 1.4 100 1.8 $\pm$ 0.3 12.4 $\pm$ 3.5 80 9.85 $\pm$ 6.66
8 0.0 $\pm$ 0.0 40 2.7 $\pm$ 1.1 7.0 $\pm$ 4.2 20 0.16 $\pm$ 0.0
9 NaN 0 NaN NaN 0 NaN
Table 6: Computation Time (sec) - greedy+SAS
$N$ $t_{\text{R101,greedy+SAS}}$ $t_{\text{R101,CQM}}$ $t_{\text{R201,greedy+SAS}}$ $t_{\text{R201,CQM}}$
5 3.29 5.09 4.12 5.11
6 6.03 5.06 9.01 5.14
7 10.32 5.07 13.64 5.21
8 19.51 5.2 25.14 5.33
9 41.1 5.3 65.05 5.45
10 51.14 5.6 93.61 5.63
15 230.58 6.77 618.08 6.74
25 21.6 10.96 43.1 10.93
50 433.16 65.88 744.42 104.67
(a) Solomon R101 Examples
(b) Solomon R201 Examples
Figure 5: Average relative optimality gap comparing greedy+SAS and CQM averaged over ten examples with standard deviation.
6.2 Results - greedy+SAS vs CQM
To show the effectiveness of our decomposition, greedy+SAS and CQM are also compared. The relative optimality gap for various problem sizes is shown in Fig. 5(b), and the computation times are shown in Table 6. At small problem sizes, the rel. The optimality gap remains low, and at large problem sizes, the optimality gap for the greedy algorithm is significantly lower than that of CQM. This trend is even more pronounced for the $R$201 examples, showing that the decomposition performs better than solving Problem (4) using a state-of-the-art hybrid solver CQM.
6.3 Results - greedy+QPU vs CQM
In Tab. 5, we can see results averaged over ten examples for solving greedy+QPU. Because the process of embedding the logical problem on the physical device is non-deterministic, the following procedure was used for each problem size: For each of the ten random examples, the embedding process was executed 10 times. The embedding using the fewest number of physical qubits was used to execute the remaining iterations of Algorithm 1. The average embedding time was recorded and added to the time taken to complete the remaining iterations of the algorithm. The statistics for optimality gap and computation time are in Tab. 5. For problem sizes $N\in\{5,6,7\}$, a valid embedding was found for all ten random examples. For $N=8$, $40\%$/$20\%$ of the examples, a valid embedding was found for $R$101/$R$201, respectively, and for $N=9$, none of the examples found a valid embedding. Given the current QPU capabilities, a hybrid solver for the QUBO subproblems is necessary for practical-scale problems. For the same examples, CQM always found the optimal solution.
6.4 Results - greedy+QPU vs QPU
To evaluate robustness to noise, greedy+QPU and QPU were compared. For a fair benchmark, the parameters of the DWaveSampler were tuned in pure QPU. An annealing pause was used by setting the schedule to [0.0, 0.0], [10.0, 0.4], [50.0, 0.4], [120, 1.0]. Annealing pauses are shown to perform well for scheduling problems [15]. Finally, to increase the chance of the DWaveSampler finding a feasible solution, both the annealing time and the number of samples were adjusted to the largest values allowed by D-Wave’s QPU access time: $50\mu s$ and $M$=2000, respectively. In greedy+QPU, we left the annealing parameters in DWaveSampler to their defaults. Even after these enhancements DWaveSampler still fails to find a feasible solution in any of the ten examples for problem sizes greater than N=5 (with/without pausing). For N=5, only 20% of the solutions were feasible (not shown in Tab. 4). With pausing [15], surprisingly, this probability decreased to 10%. However, greedy+QPU still maintains an average relative optimality gap below 33%. The total time of Alg. 1 and the average relative optimality gap are shown in Tab. 4. The average number of logical/physical qubits is also reported ($X$) to show the scaling of this formulation on the binary variables. As $N$ increases, the number of logical qubits begins to attenuate. Only one discretization value within each time window is needed. As a timetable includes more customers, more of the time windows overlap - avoiding the need to create additional time discretizations.
7 Conclusions and Future Work